OpenStudy (anonymous):

Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 (2 points): Use calories as the y−coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned

OpenStudy (anonymous):

part1: there r two pts to help u decide the slope, point A is (12,36)B is (18,54) then in order to get the slope: (y-y')/(x-x")=18/6=3 the slope is 3 which means one more minute burn 3 calories part2:u already know the slope, so y=3x+b, plug A or B in it, then u ll get it part3:when u get the func, just plug X=35 in the fct

OpenStudy (anonymous):

hope it helps~

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

srry my pc acted up, anyways i had a question about part 2

OpenStudy (anonymous):

@mymathcourses

OpenStudy (anonymous):

i didnt understand what u meant when u said plug either a or b into the equation